OpenStudy (anonymous):

f(x)=x-1 g(x)=2x^2-2x-5 Find g(f(x)) @Jhannybean

OpenStudy (jhannybean):

\[\large g(x) = 2x^2 -2x -5\]\[\large f(x) = x-1 \] now take the function of g(x) and wherever you see an "x" in that function, replace it with the "f(x)" formula \[\large g(f(x)) = 2(x-1)^2 -2(x-1) -5\]Can you solve this now? :)

OpenStudy (jhannybean):

*f(x) function* not formula, sorry about that.

OpenStudy (anonymous):

2x^2-6x-1?

OpenStudy (jhannybean):

Good job :D

OpenStudy (anonymous):

I still don't really feel %100 on these, and I have 8 more like this lol

OpenStudy (jhannybean):

Like the substitution ones?

OpenStudy (anonymous):

like the ones we just did.......

OpenStudy (jhannybean):

ah. Then just follow the same method we used here. Just for practice,could you tellme how you would find \(\large f(g(x))?\)

OpenStudy (anonymous):

multiply g*x and then multiply the result * f?

OpenStudy (jhannybean):

No we would just take the original g(x) function and input it into f(x) like so \[\large f(g(x)) = (2x^2-2x-5)-1\]

OpenStudy (anonymous):

sorry had to let the dogs out

OpenStudy (jhannybean):

no problem, lol